You are browsing a read-only backup copy of Wikitech. The live site can be found at wikitech.wikimedia.org

Nova Resource:Tools.openstack-browser/SAL

From Wikitech-static
Jump to navigation Jump to search

2021-09-26

  • 15:03 majavah: upgrading to python 3.9

2021-07-04

  • 08:35 majavah: deploying patch to include cinder usage on front page storage space counts

2021-07-03

  • 17:09 majavah: deploying patch to display list of trove databases T282801

2021-05-15

2021-05-03

  • 11:07 Majavah: update to fix cinder quota display

2021-04-29

  • 18:19 Majavah: deploy changes for T211393

2021-03-26

  • 11:32 Majavah: deploy cinder quota display

2021-03-25

2020-12-03

  • 09:17 arturo: webservice restart

2020-10-05

  • 17:44 wm-bot: <bd808> Purged cache of puppet roles for T264649

2020-09-08

2020-07-23

2020-05-06

2020-02-20

  • 11:29 arturo: restarting webservice as the page is mostly not responding (in both domains)

2020-02-14

  • 22:02 Krenair: Restarting webservice as everything was waiting and then returning 504 etc

2020-02-11

  • 10:44 arturo: restart webservice to create new ingress object for toolforge.org

2020-02-10

  • 11:03 arturo: restarted webservice by hand for nothing, given we don't have here yet the webservice version I was expecting

2020-02-06

  • 11:04 arturo: restarted webservice to test how it works with latest front proxy changes (T234617) Then realized we lack newer version of webservice and this restart was for nothing

2020-01-03

  • 22:24 bd808: Migrating to new kubernetes cluster

2019-10-09

  • 23:02 bd808: Restarting to see if the webservice ends up on a less loaded exec node

2019-09-13

  • 04:29 bd808: Updated to a468c10 (Point at new Cloud VPS puppetmaster)

2019-02-11

  • 10:59 gtirloni: Deployed change to show hypervisor in VM list

2018-11-01

  • 22:08 Krenair: forgot to log earlier: pulled commit 2beb3e1 for T164332
  • 19:28 bd808: Added arturo, gtirloni, and Krenair as tool maintainers. Removed madhu.
  • 18:11 bd808: Update to 88f9554 (Show instance hypervisors) T208502

2018-10-01

2018-08-23

  • 13:09 bstorm_: Added multi-region checks to nova lookups -- yesterday

2018-05-17

  • 14:16 chasemp: restart

2018-05-15

  • 20:10 bd808: Restarting pod and purging cached data seems to have fixed the webservice
  • 20:04 bd808: Returning 500 errors, investigating

2018-04-13

  • 21:49 bd808: Upgrade to b71a890 and restart

2017-08-31

  • 23:14 madhuvishy: Restarted webservice after fixing T174740

2017-05-24

  • 14:54 bd808: Added -quiet and stderr/out to /dev/null to hourly cache purge cron job

2017-04-21

  • 00:10 bd808: Deployed c491ace to display proxy information on project pages. T45580

2017-04-20

  • 17:57 bd808: Deployed 7abd6ee to show LDAP user count now that T162745 is fixed

2017-04-06

  • 15:47 bd808: Unbroken now, but LDAP user count is incorrect (paging limit)
  • 15:29 bd808: Testing changes and they seem to be a bit broken